How many liters of 25 percent salt solution must be added to 20 liters of a 12 percennt solution to make a solution that is 20 percent salt?
Answer by mananth(16946)

solution 12----------------20
solution 25----------------x
Total 20 ----------------20+x
12*20+25*x=20(20+x)
240+25x =400+20x
25x-20x =400-240
5x=160
/5
x=32 liters of 25%